4. (a) Supply proof of the Menelaus Theorem concerning a transversal line L cutting the sides of ΔABC at points X,Y,Z respectively. (Hint) Drop perpendicular line segments from A, B, C to L and use similar triangles b)Centuries after Menelaus, Ceva discovered the Theorem that if P,Q, R are points on BC, CA and AB respectively so that AP, BQ, CR meet at a single point K, thern AR BP co RB PC QA Prove Cevas theorem and its converse, by using Menelaus Theorem. (Hint: There are 6 triangles, each sharing a common side and a common angle with ΔABC, and each is being cut transversally by an appropriate line)
Answer #1

Let a triangle ABC with extended side BC.

Now Let a traversal line L cuts sides at X, Y and Z.

Now Let we drop perpendiculars from A, B and C to L.

From Similar triangles ---

\frac{AZ}{BZ} = \frac{AP}{BQ} ______(1)
\frac{BC}{CX} = \frac{BQ}{CR} ______(2)

\frac{AY}{CY} = \frac{CR}{AP} ______(3)

Multiply all three equations ---

\frac{AY}{CY} \times \frac{BC}{CX}\times \frac{AZ}{BZ} = \frac{CR}{AP} \times \frac{BQ}{CR} \times \frac{AP}{BQ}

\frac{AY}{CY} \times \frac{BC}{CX}\times \frac{AZ}{BZ} = 1

Hence proved.